OpenStudy (elleblythe):

Simplify [(4^8xy^9)(2-1)^10/(-2)^17(x^-1y^2)^2]^-3

OpenStudy (anonymous):

\[[\frac{ (4^{8}xy^{9})(2-1)^{10} }{ (-2)^{17}(x^{-1}y^{2})^{2}}]^{-3}\] is it like this??? @elleblythe

OpenStudy (elleblythe):

@curious9876 yes! how do you simplify that? if you could show solutions please! just wanna make sure my answer is correct

OpenStudy (anonymous):

first subtract 1 from 2 that is 1and 1^10=1

OpenStudy (anonymous):

4^8=65536

OpenStudy (anonymous):

so the numerator can be be simplified as 65536xy^9.1=65536.xy^9

OpenStudy (anonymous):

nw do the same for the denominator \[(-2)^{17}=???\]

OpenStudy (anonymous):

@elleblythe

OpenStudy (anonymous):

@elleblythe \[(-2)^{17}=???\]

OpenStudy (anonymous):

AFTER FINDING THAT u have to simplify (x^-1 y^2)^2

OpenStudy (elleblythe):

Can't we just simplify the 4^8 to make it (2^2)^8 so it becomes 2^16 then we can cancel out the exponents of 2 in the numerator and denominator?

OpenStudy (anonymous):

ok u can do in that way too if u want to do like that then first write -2^17= -2.(2)^16

OpenStudy (anonymous):

nw u can cancel 2^16 and 2^16 from the nmrtr and dnmntr

OpenStudy (anonymous):

nw our expression becomes \[[\frac{ xy^{9}}{ -2.x^{-2}.y^{4}}]^{-3}\]

OpenStudy (anonymous):

nw to make the exponent of x in the denominator positive transfer it to the numerator nw the expression becomes\[[\frac{ xy^{9}.x^{2} }{ -2.y^{4} }]^{-3}\]

OpenStudy (anonymous):

have u undrstd till here @elleblythe

OpenStudy (elleblythe):

so what's your final answer?

OpenStudy (anonymous):

nw snc x and x^2 are multiplied here and the bases are same i.e. x we have to add the exponents \[[\frac{ x^{1+2}.y^{9 }}{ -2.y^{4} }]^{-3}\]

OpenStudy (anonymous):

nw add 1 and 2 =3 and cancel y^9 and y^4 wht l u get???@elleblythe

OpenStudy (anonymous):

@elleblythe

OpenStudy (elleblythe):

is it -8/x^9y^15?? @Curious9876

OpenStudy (anonymous):